OpenStudy (anonymous):

Nadia had money in a savings account at a Savings and Loan, but the Savings and Loan closed. What act protects her money in the savings account?

OpenStudy (anonymous):

A = P(1 + r/n)^(nt) r = 0.0725 n = 3 t = 2 A = 32000 Solve for P 32000 = P(1 + 0.0725/2)^(2*3) 32000 = P(1.03625)^6 32000 = 1.23819P P = $25844.18